What is 5 3/4 inches minus 1/8 inch?
Answer by nerdybill(7384)

What is 5 3/4 inches minus 1/8 inch?
convert 3/4 to eighths:
3/4 * 2/2 = 6/8
6/8 - 1/8 = 5/8
answer: 5 and 5/8